# Can Diffusion Models Provide Rigorous Uncertainty Quantification for Bayesian Inverse Problems?

### v1

@@ Abstract
In recent years, the ascendance of diffusion modeling as a state-of-the-art generative modeling approach has spurred significant interest in their use as priors in Bayesian inverse problems. However, it is unclear how to optimally integrate a diffusion model trained on the prior distribution with a given likelihood function to obtain posterior samples. While algorithms that have been developed for this purpose can produce high-quality, diverse point estimates of the unknown parameters of interest, they are often tested on problems where the prior distribution is analytically unknown, making it difficult to assess their performance in providing rigorous uncertainty quantification. In this work, we introduce a new framework, Bayesian Inverse Problem Solvers through Diffusion Annealing (BIPSDA), for diffusion model based posterior sampling. The framework unifies several recently proposed diffusion model based posterior sampling algorithms and contains novel algorithms that can be realized through flexible combinations of design choices. Algorithms within our framework were tested on model problems with a Gaussian mixture prior and likelihood functions inspired by problems in image inpainting, x-ray tomography, and phase retrieval. In this setting, approximate ground-truth posterior samples can be obtained, enabling principled evaluation of the performance of the algorithms. The results demonstrate that BIPSDA algorithms can provide strong performance on the image inpainting and x-ray tomography based problems, while the challenging phase retrieval problem, which is difficult to sample from even when the posterior density is known, remains outside the reach of the diffusion model based samplers.
@@ 1 INTRODUCTION
Inverse problems, which aim to estimate an unknown parameter of interest from observed data (measurements), provide a principled way to integrate data with existing scientific knowledge. However, many important inverse problems throughout the sciences are ill-posed—the measurements alone do not contain enough information to uniquely and stably estimate the parameter of interest. In the Bayesian inverse problem framework, this difficulty is addressed through the integration of prior knowledge regarding the parameter of interest, which enables both point estimation of the unknown parameter and rigorous uncertainty quantification [1, 2]. This can facilitate risk-informed analysis and decision making, which is particularly important in the context of safety-critical systems (e.g., medical imaging [3] or earthquake detection [4] systems), where accurate risk assessments can save lives.
@@
The success of generative modeling [5, 6] in capturing the structure of complex high-dimensional probability distributions in recent years has spurred considerable interest in the use of generative modeling to overcome this limitation [7]. In particular, over the last four years the rise to prominence of diffusion modeling [8, 9, 10] as a state-of-the-art generative modeling approach has been coupled with significant interest in their use as prior distributions in Bayesian inference. Here a common strategy is to obtain posterior samples by pretraining a diffusion model on samples from the prior distribution and integrating the pretrained model with the likelihood function (which is assumed to be known analytically) at inference time [11, 12, 13, 14, 15]. This general strategy has the advantage of enabling the same diffusion model to be used in conjunction with many different data acquisition designs without retraining, and has shown considerable promise. However, there is no consensus on how to optimally integrate the diffusion model and the likelihood function. While several algorithms have been proposed for this purpose, test-bed problems with no analytically-known ground-truth prior are often used to illustrate the performance of such methods. In this setting, the performance of the algorithms is often assessed using sample quality metrics like peak signal-to-noise ratio, as well as heuristic arguments regarding the sample diversity. This makes it difficult to assess the accuracy of these methods in capturing the global structure of the posterior.
@@
In this paper, we introduce a unified framework, Bayesian Inverse Problem Solvers through Diffusion Annealing (BIPSDA), for solving Bayesian inverse problems with diffusion models. The framework generalizes and extends two recently proposed algorithms, Diffusion Annealing Posterior Sampling (DAPS) [16] and DiffPIR [17], for solving Bayesian inverse problems with diffusion models. These two algorithms have achieved strong performance on a number of canonical imaging reconstruction problems, including non-linear problems such as phase retrieval that were previously too difficult for diffusion model based solvers [16]. In our framework, the DAPS and DiffPIR algorithms can be recovered through specific design choices, while new algorithms can be realized through exploration of the algorithmic design space. We provide examples of several novel algorithms within our framework, including a class of approaches inspired by the randomize-then-optimize (RTO) method [18, 19] for sampling from posterior distributions.
@@
We systematically evaluated the performance of algorithms in the proposed framework, including the DAPS and DiffPIR algorithms and the novel RTO-based algorithms, on a set of model problems. Each model problem uses a Gaussian mixture prior. This choice of prior has two key advantages. First, under this choice the posterior density is known and approximate ground-truth posterior samples can be obtained, enabling rigorous analysis of the performance of the BISPDA algorithms. Second, under this choice the noisy prior scores, a key component of diffusion models that is learned from data, can be formed analytically. This allows us to decompose the error in the posterior sampler into two components: error inherent to the algorithm, and error due to incorrect modeling of the prior distribution. For the likelihood functions, we considered benchmark problems inspired by classic image restoration/reconstruction problems: simple linear inpainting problems (in both low and high noise regimes), as well as nonlinear x-ray tomography and phase retrieval based problems.
@@
In each experiment, the performance of the BIPSDA algorithms was evaluated by comparing the algorithm samples with the approximate ground-truth samples using four different metrics: the central moment discrepancy (CMD) [20], maximum mean discrepancy (MMD) [21], and the errors in both the predicted posterior mean and pointwise variance. This enables principled evaluation of the ability of various BIPSDA algorithms to capture both the local and global posterior structure.
@@
The remainder of this paper is organized as follows. In Section 2, we provide relevant background on Bayesian inverse problems and diffusion models. In Section 3 we introduce the proposed BISPDA framework, discuss its relationship to previously proposed algorithms, and provide examples of algorithms that can be realized within the framework. Section 4 provides details regarding the numerical studies, while results are shown in Section 5. A discussion of the results and the conclusion is given in Section 6.
@@ 2.2 DIFFUSION MODELS
Broadly, the goal of generative modeling can be described as follows: given a set of samples =$\mathcal{M}=\{\mathbf{m}_{i}\}_{i=1}^{N_{s}}$ from a distribution of interest with density function $\pi_{0}(\mathbf{m})$, obtain a set of samples from $\pi_{0}(\mathbf{m})$ that are not in $\mathcal{M}$. In diffusion modeling, these samples are obtained by reversing a predefined noising process [9, 10]. This noising process can be written as a stochastic differential equation (SDE). In particular, the variance exploding (VE) form of the diffusion SDE can be written as follows [10]:
@@ 2.3 HIJACKING APPROACHES
The ascendance of diffusion models as a state-of-the-art generative modeling technique has spurred significant research on their use in the context of Bayesian inverse problems. In particular, there has been substantial interest in leveraging diffusion models trained on the prior distribution of a given inverse problem to sample from the posterior distribution, i.e., setting $\pi_{0}=\pi_{\mathrm{pr}}$. This general strategy has the advantage of enabling the same diffusion model to be used to sample from many posterior distributions corresponding to different choices of likelihood function without retraining the score model. In the remainder of this section, we discuss one algorithmic framework, which we refer to as the hijacking class of approaches, that follows this general strategy.
@@
[…]
@@ 3.1 THE FRAMEWORK
Concretely, given the current iterate $\mathbf{m}(t)$ and the measurements $\mathbf{y}$, the goal of the prediction stage is to obtain an approximate sample $\mathbf{m}(0)$ from the prediction distribution $\pi_{0\mid t,\mathbf{y}}(\mathbf{m}(0)\mid\mathbf{m}(t),\mathbf{y})$. In the corruption stage, $\mathbf{m}(t-\Delta t)$ is obtained by adding noise back to $\mathbf{m}(0)$, i.e., by sampling from the Gaussian noising distribution $\pi_{t-\Delta t\mid 0}(\mathbf{m}(t-\Delta t)\mid\mathbf{m}(0))$. In [16], Zhang et al prove that if $\mathbf{m}(t)$ is a sample from $\pi_{t\mid\mathbf{y}}(\mathbf{m}(t)\mid\mathbf{y})$, then this procedure yields a sample from $\pi_{t-\Delta t\mid\mathbf{y}}(\mathbf{m}(t-\Delta t)\mid\mathbf{y})$. For sufficiently large $T$, once can assume that $\pi_{T\mid\mathbf{y}}(\mathbf{m}(T)\mid\mathbf{y})\approx\mathcal{N}(\mathbf{m}(T);\mathbf{0},\sigma^{2}(T)\mathbf{I})$. So starting from $\mathbf{m}(T)$, the prediction and corruption steps can be iteratively applied with the timestep annealed from $T$ down to 00 to yield a sample from the posterior distribution.
@@
1:  Input: Decreasing timesteps $[t_{N_{A}},t_{N_{A}-1},\cdots,t_{0}]$, likelihood function $\pi_{\mathrm{like}}$, noise schedule $\sigma(t)$, trained score model $s_{\boldsymbol{\theta}^{*}}(\mathbf{m}(t),t)$
@@ 3.2 RELATIONSHIP TO PRIOR WORK
Over the last two years, several diffusion-based Bayesian inverse problem solvers have been proposed that use decoupled noise annealing. Of these approaches, the work of Zhang et al [16] is the most closely related to the proposed BIPSDA framework. In their approach, referred to as Decoupled Annealing Posterior Sampling (DAPS), the approximation of the denoising distribution (line 5 in Algorithm 1) takes the form of a Gaussian distribution, i.e., we have that
@@
Here $\mathbf{m}_{\mathrm{aprx}}$ and $\mathbf{C}_{\mathrm{aprx}}$ are the mean and covariance, respectively, of the approximation distribution. In particular, in the DAPS approach the covariance matrix is set as $\mathbf{C}_{\mathrm{aprx}}=\beta(t)^{2}\mathbf{I}$, with the noise level $\beta(t)$ chosen using heuristics. The mean of $\mathbf{m}_{\mathrm{aprx}}$ is set as an estimate of $\mathbb{E}_{0\mid t}[\mathbf{m}(0)\mid\mathbf{m}(t)]$, with the estimate obtained by solving the probability flow ODE in (5) with initial value $\mathbf{m}(t)$. Note that under a particular discretization of the probability flow ODE, this estimate corresponds with Tweedie’s formula for computing the mean and is thus exact up to error in the score model. However, in general, the probability flow ODE yields a biased estimate of the mean; see Appendix 7 for an analysis. To sample from the corresponding approximate prediction distribution (line 6 in Algorithm 1), the DAPS approach uses the Euler–Maruyama method for discretizing Langevin dynamics (without Metropolis adjustment).
Another approach that is closely related to the present work is the DiffPIR algorithm of Zhu et al [17]. Like the DAPS algorithm, this approach models the denoising distribution as a Gaussian, with the mean set as an estimate of the conditional mean $\mathbb{E}_{0\mid t}[\mathbf{m}(0)\mid\mathbf{m}(t)]$ and the covariance matrix chosen to be a scalar multiple of the identity. However, in their approach, the conditional mean is estimated using Tweedie’s formula, which, as noted above, is exact up to error in the score model. Further, the noise level of the covariance matrix is set so that $\mathbf{C}_{\mathrm{aprx}}=\sigma^{2}(t)\mathbf{I}$, so unlike DAPS the covariance matrix does not depend on a user-specified hyperparameter.
While the DAPS and DiffPIR algorithms both use Gaussians to model the approximation distribution, the DiffPIR algorithm does not use Langevin dynamics to sample from the corresponding prediction distribution. Instead, inspired by plug-and-play algorithms for image restoration (see, e.g., [31]), the DiffPIR algorithm obtains a “sample” by solving the maximum a posteriori (MAP) estimation problem corresponding to the prediction distribution. While lacking theoretical guarantees, it is empirically observed that DiffPIR can produce very diverse samples (see Figure 5 in [17]). Furthermore, the approach enables the use of efficient, problem-specific MAP solvers. For example, when the measurement noise is additive Gaussian and the denoising distribution is modeled as a Gaussian, the MAP point is given by evaluation of the proximal operator corresponding to the forward model $f(\cdot)$, for which specialized algorithms are available for many classes of forward models [32].
Another difference between DiffPIR and DAPS algorithms lies in the “corruption” stage of the algorithms (line 7 of Algorithm 1). In DAPS, as in the BIPSDA framework, the corruption stage corresponds to sampling from the Gaussian noising distribution. In DiffPIR, additional flexibility is enabled through the use of the denoising diffusion implicit models (DDIM) technique introduced by Song et al [33]. In particular, in this approach, $\mathbf{m}(t_{i-1})$ depends explicitly on both $\mathbf{m}(0)$ and $\mathbf{m}(t_{i})$, with an additional hyperparameter introduced to control the weighting of the two terms. For a particular choice of this hyperparameter, the explicit dependence of $\mathbf{m}(t_{i-1})$ on $\mathbf{m}(t_{i})$ disappears, and the corruption approach of DAPS and the proposed BIPSDA framework is recovered.
3.3 EXAMPLE ALGORITHMS
Regarding the approximation of the denoising distribution, we first note that in principle, arbitrarily complex approximations of the denoising distribution can be realized by first sampling from the denoising distribution $\pi_{0\mid t}(\mathbf{m}(0)\mid\mathbf{m}(t))$ using the reverse SDE and fitting a model to the samples. However, these approximations may be computationally prohibitive in practice, as they would require many solves of the reverse SDE and a model-fitting step at each BIPDSA iteration. We therefore restrict our focus to Gaussian approximations that do not require fitting a model to samples. In particular, we consider three different methods for setting the mean $\mathbf{m}_{\mathrm{aprx}}$ and covariance $\mathbf{C}_{\mathrm{aprx}}$ of the Gaussian approximation in (9). In the first method, which is based on the DAPS algorithm, the mean is obtained by solving the probability flow ODE, and the covariance is chosen as $\mathbf{C}_{\mathrm{aprx}}=\beta(t)^{2}\mathbf{I}$, where here we set $\beta(t)=\sigma(t)$ for simplicity. In the second method, the mean and covariance are set as in the DiffPIR algorithm, with $\mathbf{m}_{\mathrm{aprx}}$ obtained using Tweedie’s formula and $\mathbf{C}_{\mathrm{aprx}}=\sigma(t)^{2}\mathbf{I}$. Finally, in the third method, which is inspired by the hijacking approach of Boys et al [26], both the mean and covariance of the Gaussian distribution are set using the generalized version of Tweedie’s formula [30]. In the remainder of this paper, we refer to the three methods described above respectively as the ODE, Tweedie Uncorrelated (TU), and Tweedie Correlated (TC) methods for approximating the denoising distribution.

### v2

# Benchmarking Diffusion Annealing-Based Bayesian Inverse Problem Solvers

@@ Abstract
In recent years, the ascendance of diffusion modeling as a state-of-the-art generative modeling approach has spurred significant interest in their use as priors in Bayesian inverse problems. However, it is unclear how to optimally integrate a diffusion model trained on the prior distribution with a given likelihood function to obtain posterior samples. While algorithms developed for this purpose can produce high-quality, diverse point estimates of the unknown parameters of interest, they are often tested on problems where the prior distribution is analytically unknown, making it difficult to assess their performance in providing rigorous uncertainty quantification. Motivated by this challenge, this work introduces three benchmark problems for evaluating the performance of diffusion model based samplers. The benchmark problems, which are inspired by problems in image inpainting, x-ray tomography, and phase retrieval, have a posterior density that is analytically known. In this setting, approximate ground-truth posterior samples can be obtained, enabling principled evaluation of the performance of posterior sampling algorithms. This work also introduces a general framework for diffusion model based posterior sampling, Bayesian Inverse Problem Solvers through Diffusion Annealing (BIPSDA). This framework unifies several recently proposed diffusion-model-based posterior sampling algorithms and contains novel algorithms that can be realized through flexible combinations of design choices. We tested the performance of a set of BIPSDA algorithms, including previously proposed state-of-the-art approaches, on the proposed benchmark problems. The results provide insight into the strengths and limitations of existing diffusion-model based posterior samplers, while the benchmark problems provide a testing ground for future algorithmic developments.
@@ 1 INTRODUCTION
Inverse problems, which aim to estimate an unknown parameter of interest from observed data (measurements), provide a principled way to integrate data with existing scientific knowledge [1, 2]. However, many important inverse problems throughout the sciences are ill-posed [3, 4]—the measurements alone do not contain enough information to uniquely and stably estimate the parameter of interest [5]. In the Bayesian inverse problem framework, this difficulty is addressed through the integration of prior knowledge regarding the parameter of interest, which enables both point estimation of the unknown parameter and rigorous uncertainty quantification [6, 7]. This can facilitate risk-informed analysis and decision making, which is particularly important in the context of safety-critical systems (e.g., medical imaging [8] or earthquake detection [9] systems), where accurate risk assessments can save lives.
@@
The success of generative modeling [10, 11] in capturing the structure of complex high-dimensional probability distributions in recent years has spurred considerable interest in the use of generative modeling to overcome this limitation [12]. In particular, over the last four years the rise to prominence of diffusion modeling [13, 14, 15] as a state-of-the-art generative modeling approach has been coupled with significant interest in their use as prior distributions in Bayesian inference. Here a common strategy is to obtain posterior samples by pretraining a diffusion model on samples from the prior distribution and integrating the pretrained model with the likelihood function (which is assumed to be known analytically) at inference time [16, 17, 18, 19, 20, 21]. This general strategy has the advantage of enabling the same diffusion model to be used in conjunction with many different data acquisition designs without retraining, and has shown considerable promise. However, there is no consensus on how to optimally integrate the diffusion model and the likelihood function. While several algorithms have been proposed for this purpose, test-bed problems with no analytically-known ground-truth prior are often used to illustrate the performance of such methods. In this setting, the performance of the algorithms is often assessed using sample quality metrics like peak signal-to-noise ratio, as well as heuristic arguments regarding the sample diversity. This makes it difficult to assess the accuracy of these methods in capturing the global structure of the posterior.
@@
In this paper, we provide a general framework, Bayesian Inverse Problem Solvers through Diffusion Annealing (BIPSDA), which—by abstracting the algorithmic components of the Diffusion Annealing Posterior Sampling (DAPS) method described in [22]—provides an unified formulation for developing and analyzing annealing-based solvers for Bayesian inverse problems with priors implicitly defined by a diffusion model. The framework generalizes and extends two recently proposed algorithms, DAPS [22] and DiffPIR [23], for solving Bayesian inverse problems with diffusion models. These two algorithms have achieved strong performance on a number of canonical imaging reconstruction problems, including non-linear problems such as phase retrieval that were previously considered too difficult for diffusion-model-based solvers [22]. In the unified framework, the DAPS and DiffPIR algorithms can be recovered through specific design choices, while new algorithms can be unveiled through exploration of the rich algorithmic design space. An original contribution of our work is the use of randomize-then-optimize (RTO) techniques, originally proposed in [24, 25, 26] for approximate sampling from posterior distributions, to solve a sampling subproblem that arises in the BIPSDA framework. Here the RTO technique transforms this subproblem into an optimization problem (as in DiffPIR) that can be solved with “off-the-shelf” deterministic numerical optimization methods, while still accurately accounting for measurement noise statistics (as in the DAPS method).
@@
We systematically evaluated the performance of algorithms in the proposed framework, including the DAPS and DiffPIR algorithms and the novel RTO-based algorithms, on a set of model problems. Each model problem uses a Gaussian mixture prior. This choice of prior has two key advantages. First, under this choice the posterior density is known and approximate ground-truth posterior samples can be obtained, enabling rigorous analysis of the performance of the BIPSDA algorithms. Second, under this choice the noisy prior scores, a key component of diffusion models that is learned from data, can be formed analytically. This allows us to decompose the error in the posterior sampler into two components: error inherent to the algorithm, and error due to incorrect modeling of the prior distribution. For the likelihood functions, we considered benchmark problems inspired by classic image restoration/reconstruction problems: simple linear inpainting problems (in both low and high noise regimes), as well as nonlinear x-ray tomography and phase retrieval based problems.
@@
In each experiment, the performance of the BIPSDA algorithms was evaluated by comparing the samples produced by each algorithm with the approximate ground-truth samples using four different metrics: the central moment discrepancy (CMD) [27], maximum mean discrepancy (MMD) [28], and the errors in both the predicted posterior mean and pointwise variance. This enables principled evaluation of the ability of various BIPSDA algorithms to capture both the local and global posterior structure.
@@
The remainder of this paper is organized as follows. In Section 2, we provide relevant background on Bayesian inverse problems and diffusion models. In Section 3 we introduce the proposed BIPSDA framework, discuss its relationship to previously proposed algorithms, and provide examples of algorithms that can be realized within the framework. Section 4 provides details regarding the numerical studies, while results are shown in Section 5. A discussion of the results and the conclusion is given in Section 6.
@@ 2.2 DIFFUSION MODELS
Broadly, the goal of generative modeling can be described as follows: given a set of samples $\mathcal{M}=\{\mathbf{m}_{i}\}_{i=1}^{N_{s}}$ from a distribution of interest with density function $\pi_{0}(\mathbf{m})$, obtain a set of samples from $\pi_{0}(\mathbf{m})$ that are not in $\mathcal{M}$. In diffusion modeling, these samples are obtained by reversing a predefined noising process [14, 15]. This noising process can be written as a stochastic differential equation (SDE). In particular, the variance exploding (VE) form of the diffusion SDE can be written as follows [15]:
@@ 2.3 HIJACKING APPROACHES
The ascendance of diffusion models as a state-of-the-art generative modeling technique has spurred significant research on their use in the context of Bayesian inverse problems. In particular, there has been substantial interest in leveraging diffusion models trained on the prior distribution of a given inverse problem to sample from the posterior distribution, i.e., setting $\pi_{0}=\pi_{\mathrm{pr}}$; see [21] for a survey of these approaches. This general strategy has the advantage of enabling the same diffusion model to be used to sample from many posterior distributions corresponding to different choices of likelihood function without retraining the score model. In the remainder of this section, we discuss one algorithmic framework, which we refer to as the hijacking class of approaches, [^1] These approaches are referred to as “Explicit Approximations for Measurement Matching” approaches in [21]. that follows this general strategy.
@@
$\quad=\int\pi_{\mathrm{like}}(\mathbf{y}\mid\mathbf{m}(0))\,\pi_{0\mid t}(\mathbf{m}(0)\mid\mathbf{m}(t))\;d\mathbf{m}(0),$ (8)
@@ 3.1 THE FRAMEWORK
Concretely, given the current iterate $\mathbf{m}(t)$ and the measurements $\mathbf{y}$, the goal of the prediction stage is to obtain an approximate sample $\mathbf{m}(0)$ from the prediction distribution $\pi_{0\mid t,\mathbf{y}}(\mathbf{m}(0)\mid\mathbf{m}(t),\mathbf{y})$. In the corruption stage, $\mathbf{m}(t-\Delta t)$ is obtained by adding noise back to $\mathbf{m}(0)$, i.e., by sampling from the Gaussian noising distribution $\pi_{t-\Delta t\mid 0}(\mathbf{m}(t-\Delta t)\mid\mathbf{m}(0))$. In [22], Zhang et al prove that if $\mathbf{m}(t)$ is a sample from $\pi_{t\mid\mathbf{y}}(\mathbf{m}(t)\mid\mathbf{y})$, then this procedure yields a sample from $\pi_{t-\Delta t\mid\mathbf{y}}(\mathbf{m}(t-\Delta t)\mid\mathbf{y})$. For sufficiently large $T$, once can assume that $\pi_{T\mid\mathbf{y}}(\mathbf{m}(T)\mid\mathbf{y})\approx\mathcal{N}(\mathbf{m}(T);\mathbf{0},\sigma^{2}(T)\mathbf{I})$. So starting from $\mathbf{m}(T)$, the prediction and corruption steps can be iteratively applied with the timestep annealed from $T$ down to 0 to yield a sample from the posterior distribution.
@@
1:Input: Decreasing timesteps $t_{N_{A}},t_{N_{A}-1},\cdots,t_{0}]$, likelihood function $\pi_{\mathrm{like}}$, noise schedule $\sigma(t)$, trained score model $s_{\boldsymbol{\theta}^{*}}(\mathbf{m}(t),t)$
@@ 3.2 RELATIONSHIP TO PRIOR WORK
Over the last two years, several diffusion-based Bayesian inverse problem solvers have been proposed that use decoupled noise annealing. Of these approaches, the work of Zhang et al [22] is the most closely related to the proposed BIPSDA framework. In their approach, referred to as Decoupled Annealing Posterior Sampling (DAPS), the approximation of the denoising distribution (line 5 in Algorithm 1) takes the form of a Gaussian distribution with mean $\mathbf{m}_{\mathrm{aprx}}$ and covariance $\mathbf{C}_{\mathrm{aprx}}=\beta(t)^{2}\mathbf{I}$, where the noise level $\beta(t)$ is chosen using heuristics. The mean of $\mathbf{m}_{\mathrm{aprx}}$ is set as an estimate of $\mathbb{E}_{0\mid t}[\mathbf{m}(0)\mid\mathbf{m}(t)]$, with the estimate obtained by solving the probability flow ODE in (5) with initial value $\mathbf{m}(t)$. To sample from the corresponding approximate prediction distribution (line 6 in Algorithm 1), the DAPS approach uses MCMC algorithms such as the Euler–Maruyama method for discretizing Langevin dynamics (without Metropolis adjustment), or the Hamiltonian Monte Carlo algorithm [38].
Another approach that is closely related to the present work is the DiffPIR algorithm of Zhu et al [23]. Like the DAPS algorithm, this approach models the denoising distribution as a Gaussian, with the mean set as an estimate of the conditional mean $\mathbb{E}_{0\mid t}[\mathbf{m}(0)\mid\mathbf{m}(t)]$ and the covariance matrix chosen to be a scalar multiple of the identity. However, unlike DAPS, DiffPIR estimates the conditional mean using Tweedie’s formula, which is exact up to error in the score model. Another key difference between DAPS and DiffPIR is that DiffPIR does not use Langevin dynamics to sample from the corresponding prediction distribution. Instead, inspired by plug-and-play algorithms for image restoration (see, e.g., [39]), it solves a maximum a posteriori (MAP) estimation problem corresponding to the prediction distribution. This problem can be efficiently solved using fast numerical optimization methods or, in some cases, close-form proximal operators [40]. While lacking theoretical guarantees, it is empirically observed that DiffPIR can produce very diverse samples (see Figure 5 in [23]).
3.3 SPECIFIC VARIANTS OF THE FRAMEWORK
3.3.1 Approximation of the denoising distribution
In this work, we consider three different approaches to build a Gaussian approximation to the denoising distribution (c.f. line 5 in Algorithm 1). That is, we consider approximations to $\pi_{0\mid t}(\mathbf{m}(0)|\mathbf{m}(t))$ of the form
@@ 3.3.1 Approximation of the denoising distribution
The rationale for restricting our focus to a Gaussian approximation is that, while in principle arbitrarily complex approximations of the denoising distribution can be realized by first sampling from the denoising distribution $\pi_{0\mid t}(\mathbf{m}(0)\mid\mathbf{m}(t))$ using the reverse SDE and then fitting a model to the samples, these approximations may be computationally prohibitive in practice. The approaches considered are denoted as ‘ODE’, Tweedie Uncorrelated (‘TU’), and Tweedie Correlated (‘TC’).
In the ‘ODE’ approach, which is based on the DAPS algorithm, the mean $\mathbf{m}_{\mathrm{aprx}}$ is obtained by solving the probability flow ODE in Eq. (5), and the covariance is chosen as $\mathbf{C}_{\mathrm{aprx}}=\beta(t)^{2}\mathbf{I}$, with $\beta(t)=\mathcal{O}(\sigma(t))$. User-provided hyperparameters of this approach are the marginal variance of the denoising distribution $\beta(t)^{2}$ and the schedule of time steps in the discretization of the probability flow ODE.
In the ‘TU’ approach, which is based on the DiffPIR algorithm, the mean is obtained using Tweedie’s formula, i.e.,
$\mathbf{m}_{\mathrm{aprx}}=\mathbf{m}(t)+\sigma^{2}(t)s_{\boldsymbol{\theta}^{*}}(\mathbf{m}(t),t),$
and the covariance is set as $\mathbf{C}_{\mathrm{aprx}}=\beta(t)^{2}\mathbf{I}$, with $\beta(t)=\mathcal{O}(\sigma(t))$ the only user-provided hyperparameter. We note that, as shown in Appendix 7, under a particular choice of parametrization and discretization, the probability flow ODE estimate used in the ‘ODE’ approach can be made to coincide with Tweedie formula.
Finally, in the ‘TC’ approach, which borrows some ideas from the hijacking approach of Boys et al [33], the mean $\mathbf{m}_{\mathrm{aprx}}$ is set as in the ‘TU’ approach, while the covariance is set using the generalized version of Tweedie’s formula [37], i.e.,
$\mathbf{C}_{\mathrm{aprx}}=\sigma^{2}(t)\left[\mathbf{I}+\sigma^{2}(t)\nabla_{\mathbf{m}(t)}s_{\boldsymbol{\theta}^{*}}(\mathbf{m}(t),t)\right].$
This variant is free from user-provided hyperparameters, but requires access to accurate estimates of the Jacobian of the noisy prior score.
3.3.2 Sampling from the prediction distribution
The BIPSDA framework also provides considerable flexibility regarding the choice of the sampling scheme. Three variants, denoted as Langevin (‘Lang’), maximum a posteriori estimation (‘MAP’), and randomize-then-optimize (‘RTO’), are considered here.
The ‘Lang’ variant is based on DAPS [22] and employs Markov chain Monte Carlo (MCMC) algorithms, such as Langevin dynamics or Hamiltonian Monte Carlo [38], to sample the prediction distribution. When using uncorrected Langevin dynamics the step-size and the number of time steps are key user-provided hyperparameters that control the trade-off between accuracy of the samples and computational costs.
The ‘MAP’ variant is based on DiffPIR [23] and simply computes the maximum a posteriori (MAP) point of the prediction distribution using a deterministic optimization algorithm, rather than drawing samples. While lacking solid theoretical foundation, this approach is computationally efficient. User-provided hyperparameters depends on the specific choice of the optimizer used, and, as long as the optimization problem is solved with sufficient accuracy, have limited effect on the quality of the solution.